A very noisy chain saw operated by a tree surgeon emits a total acoustic power of 32.0 W uniformly in all directions.

A) At what distance from the source is the sound level equal to 106 dB ?

B) At what distance from the source is the sound level equal to 54.0 dB ?
Answer #1

Given,

P = 32 W ;

A) Let R be the distance at which intensity I = 106 dB = 0.0398 W/m2 [Using I(dB = 10 log10 (I/Io)]

We know from the defination of intensity:

Intensity = I = Power/Area = P/A

A = P/I => 4 pi R2 = P/I

4 x 3.14 x R2 = 32 W/0.0398 W/m2

R = 8 m

Hence, distance = R = 8 m.

B)I' = 54 dB = 2.512 x 10-7 W/m2

Again using the same method:

Intensity = I = Power/Area = P/A

A = P/I => 4 pi R2 = P/I

4 x 3.14 x R2 = 32 W/ 2.512 x 10-7 W/m2

R = 3184.7 m

Hence, distance = R = 3184.7 m.